a rectangular parking lot is .045 square miles the length is .15 longer than the width. find the width and the length
Answer by lwsshak3(11628)

a rectangular parking lot is .045 square miles the length is .15 longer than the width. find the width and the length.
***
let x=width
x+.15=length
Area=width*length
.045=x(x+.15)=x^2+.15x
x^2+.15x-.045=0
(x-.15)(x+.30)=0
x=-.30 (reject)
or
x=.15
x+.15=.30
width=0.15 mi
length=0.30 mi